OpenStudy (anonymous):

Solve for x. |2x + 6| - 4 = 20 x = 4 and x = 0 x = -15 and x = 9 x = -4 and x = -2 x = -3 and x = 1 no solutions

OpenStudy (anonymous):

Firstly solve for : 2x + 6 -4 = 20 Find x from here..

OpenStudy (jiteshmeghwal9):

2x+6-4=20 2x+2=20 2x=20-2 2x=18 Can u solve it from here???

OpenStudy (anonymous):

But you have to do more.. Next you have to solve for: -(2x + 6) - 4 = 20 Or, 2x + 6 + 4 = -20 2x = -20 - 10 2x = -30 find x from here also..

OpenStudy (anonymous):

no solution?